if p=1.01atm moles=.00831 and temp is 25 celcius what is p
1
Expert's answer
2012-12-24T10:32:46-0500
You have P аnd it is 1.01 atm
For answering this question you need next equation :
PV = nRT

P-pressure
V-volume
n-amount
R-constant (8.31)
T-temperature

P = nRT/V
n = 0.00831
V = const
T = 273+25 = 298 K